OpenStudy (anonymous):

Dennis drew the line of best fit on the scatter plot shown below:

OpenStudy (anonymous):

This is the graph

OpenStudy (anonymous):

What is the approximate equation of this line of best fit in slope-intercept form? y = −15x + 2/3 y = -2 over 3x + 15 y = −15x + 3/2 y = -3 over 2x + 15